Draw the structure of an alkane that has 5 or fewer carbons and 9 primary hydrogens.
Answer #1

Primary \(\left(1^{\circ}\right)\) hydrogens mean \(\mathrm{H}\) atoms connected to \(\mathrm{C}\) atom that is bonded to only one carbon. Secondary \(\left(2^{0}\right)\) hydrogens mean \(\mathrm{H}\) atoms connected to \(\mathrm{C}\) atom that is bonded to two carbons. Tertiary \(\left(3^{\circ}\right)\) hydrogens mean \(H\) atoms connected to \(C\) atom that is bonded to three carbons. One possibility is isobutane with 9 primary hydrogens and one tertiary hydrogen. Other possibility is isopentane with 9 primary hydrogens, two secondary hydrogens and one tertiary hydrogen.

image.png